OpenStudy (anonymous):

Missed the day of class talking about T1 and T2 in equilibrium. How does one find the value of T1 in this case?

OpenStudy (anonymous):



OpenStudy (shane_b):

If they are in equilibrium then the tension forces are balanced:\[85N=T_1cos(35)\]\[T_1=\frac{85N}{cos(35)}\]

OpenStudy (anonymous):

Oh, that actually makes a lot of sense, thanks! But what if there are more than one vector?

OpenStudy (anonymous):

Then you first add the two vectors( eg. T1 and T2) and you get the sum, let it be T12. Then you want that last vector to be of the same size and opposite direction with the T3=300N vector. If they are in equilibrium.

OpenStudy (shane_b):

Here you have 2 strings supporting a 300N load and the system is in equilibrium. The tensions in the x direction must sum to 0: \[\large T_{1x}+T_{2x}=0\]\[\large T_1cos(27)+T_2cos(43)=0N\] The tensions in the y direction must sum to 300N since gravity is applying a strictly downward force:\[\large T_{1y}+T_{2y}-300N=0\]\[\large T_1sin(27)+T_2sin(43)-300N=0N\] Now you just have to solve for T1 and T2 :)